OpenStudy (anonymous):

find all subgroups and normal subgroups of group Z2 x Z2 x Z2

OpenStudy (anonymous):

any guesses?

OpenStudy (anonymous):

for one thing, you only have to find the subgroups, since \(\mathbb{Z_2^3}\) is abelian, and all subgroups must be normal

OpenStudy (anonymous):

so up to isomorphism the the subgroups will be \[\mathbb{Z_2\times Z_2}\] and \[\mathbb{Z_2}\] plus i guess the trivial ones , the group itself and zero

OpenStudy (anonymous):

i know if the group is abelian, and all subgroups are normal.. But from this link http://mathworld.wolfram.com/FiniteGroup.html , there are 16 number subgroups of group Z2 x Z2 x Z2 , but it is written only 4 number of normal. Is this true or not?

OpenStudy (anonymous):

i think since it is abelian, so all subgroups are normal.. Thank you..:)